OpenStudy (anonymous):

Find the volume of the solid that results when the region bounded by y=0, x=0, x=1, and y=x^2+1 is revolved around the y-axis using discs/washers.

OpenStudy (irishboy123):

"...using discs/washers" it's way easier just using calculus have you had a go?

OpenStudy (anonymous):

I know how to use disks/washers for most problems, but after I graphed this one, I couldn't figure it the equation. It's wrong, but I did this : \[\pi \int\limits_{0}^{2}(\sqrt{y-1}^2-(0)^2dy\]

zepdrix (zepdrix):

Hmm :) Ya if we look at the graph...

zepdrix (zepdrix):

Notice that down here, your left boundary is NOT the function x!